Px = $2, Py = $1, and income of $15
1
Expert's answer
2018-02-23T09:25:07-0500
1) So, we have that Px = $2, Py = $1, and income of $15, then we get that
$2*Qx + $1 * Qy = $15, also MCx/MCY = $2/$1, so MCx =2* MCy , then also
Qx = 2*Qy.
2) $2*2*Qy + $1 * Qy = $15, wet get QY=3, Qx = 6.